\begin{thebibliography}{99}
\bibitem{AiWa} M.~Aizenman and S.~Warzel.
\newblock {\it Random Operators: disorder effects on Quantum Spectra and Dynamics.}
\newblock Graduate Studies in Mathematics, 168. American Mathematical Society, Providence, RI, 2015. xiv+326 pp.

%\bibitem{AleksandrovPeller} A.~Aleksandrov and V.~Peller.
%\newblock Operator Lipschitz functions.
%\newblock {\it Uspekhi Mat. Nauk} 71 (2016), no. 4(430), 3--106; 
%\newblock translation in {\it Russian Math. Surveys} 71 (2016), no. 4, 605--702

%\bibitem{APPS} A.~Aleksandrov, V.~Peller, D.~Potapov and F.~Sukochev.
%\newblock Functions of normal operators under perturbations.
%\newblock {\it Adv. Math.} 226 (2011), no. 6, 5216--5251. 

%\bibitem{ABF} J.~Arazy, T.~Barton and Y.~Friedman.
%\newblock Operator differentiable functions.
%\newblock {\it Integral Equations Operator Theory} 13 (1990), no. 4, 462--487. 

%\bibitem{ABHN} W.~Arendt, C.~Batty, M.~Hieber, F.~Neubrander.
%\newblock {\it Vector-valued Laplace transforms and Cauchy problems.}
%\newblock Monographs in Mathematics, 96. Birkhäuser Verlag, Basel, 2001. xii+523 pp.
%
%\bibitem{ACS} P.~Ayre, M.~Cowling and F.~Sukochev.
%\newblock Operator Lipschitz estimates in the unitary setting.
%\newblock {\it Proc. Amer. Math. Soc.} 144 (2016), no. 3, 1053--1057.

%\bibitem{AzCaSu} N. Azamov, A. Carey, F. Sukochev.
%\newblock {\it The spectral shift function and spectral flow.}
%\newblock Comm. Math. Phys., 276 (2007), 51--91.  

\bibitem{AMSZ} N. Azamov, E. McDonald, F. Sukochev, D. Zanin.
\newblock {\it A Dixmier trace formula for the density of states,}
\newblock Commun. Math. Phys. 377, 2597–2628 (2020).


%\bibitem{AzSu2008} N. Azamov, F. Sukochev.
%\newblock {\it Spectral averaging for trace compatible operators.}
%\newblock Proc. AMS, 136 (2008), 1769--1778.  

%\bibitem{Beghin_Orsingher} L.~Beghin and E.~Orsingher.
%\newblock On the maximum of the generalized Brownian bridge.
%\newblock {\it Lithuanian Math. J.} {\bf 39} (1999), no. 2, 157--167.
%
\bibitem{Beresin_Shubin} F.\,A.\, Berezin, M.\,A.\,Shubin. 
\newblock {\it The Schr\"odinger\ equation.}
\newblock Dordrecht; Boston: Kluwer Academic Publishers, 1991

%\bibitem{BirmanSolomyak1975} M.~Birman and M.~Solomyak.
%\newblock {\it Remarks on the spectral shift function.}
%\newblock J.~Soviet Math. 3 (1975), 408-419.

%\bibitem{BirmanSolomyak1989} M.~Birman and M.~Solomyak.
%\newblock {\it Operator integration, perturbations and commutators.}
%\newblock Zap. Nauchn. Sem. Leningrad. Otdel. Mat. Inst. Steklov. (LOMI) 170 (1989), Issled. Lineĭn. Oper. Teorii Funktsiĭ. 17, 34--66, 321; 
%\newblock translation in J. Soviet Math. 63 (1993), no. 2, 129--148.

%\bibitem{BirmanSolomyakSurvey} M.~Birman and M.~Solomyak.
%\newblock {\it Double operator integrals in a Hilbert space.}
%\newblock Integral Equations Operator Theory 47 (2003), no. 2, 131--168.

%\bibitem{BourgainKlein} J.~Bourgain and  A.~Klein.
%\newblock Bounds on the density of states for Schrödinger operators.
%\newblock {\it Invent. Math.} 194 (2013), no. 1, 41--72. 
%
%\bibitem{Bratelli-Robinson-II} O.~Bratteli and D.W.~Robinson.
%\newblock {\it Operator algebras and quantum-statistical mechanics. II. Equilibrium states. Models in quantum-statistical mechanics.}
%\newblock Texts and Monographs in Physics. Springer-Verlag, New York-Berlin, 1981. xiii+505 pp.
%
%\bibitem{Brislawn-trace-class-1988} C.~Brislawn.
%\newblock Kernels of trace class operators.
%\newblock {\it Proc. Amer. Math. Soc.} 104 (1988), no. 4, 1181–1190. 
%
\bibitem{CarmonaLacroix} R.~Carmona and J.~Lacroix.
\newblock {\it Spectral theory of random Schrödinger operators.}
\newblock Probability and its Applications. Birkhäuser Boston, Inc., Boston, MA, 1990. xxvi+587 pp.

%\bibitem{action} A.~Connes.
%\newblock The action functional in noncommutative geometry.
%\newblock {\it Comm. Math. Phys.} {\bf 117} (1988), no. 4, 673--683.
%
%\bibitem{Connes} A.~Connes.
%\newblock {\it Noncommutative geometry.}
%\newblock Academic Press, Inc., San Diego, CA, 1994. xiv+661 pp.

%\bibitem{CLMSZ} A.~Connes, G.~Levitina, E.~McDonald, F.~Sukochev and D.~Zanin.
%\newblock Noncommutative geometry for symmetric non-self-adjoint operators.
%\newblock {\it J. Funct. Anal.} {\bf 277} (2019), no. 3, 889--936.

%\bibitem{CSZ} A.~Connes, F.~Sukochev and D.~Zanin.
%\newblock Trace theorem for quasi-Fuchsian groups.
%\newblock {\it Sb. Math.} {\bf 208} (2017), no. 10, 1473--1502.
%
%\bibitem{dPSW} B.~de Pagter, H.~Witvliet and F.~Sukochev.
%\newblock Double operator integrals.
%\newblock {\it J. Funct. Anal.} {\bf 192} (2002), no. 1, 52--111. 

\bibitem{DIT2001} S.~Doi, A.~Iwatsuka and T.~Mine.
\newblock The uniqueness of the integrated density of states for the Schrödinger operators with magnetic fields.
\newblock {\it Math. Z.} {\bf 237} (2001), no. 2, 335--371. 
%
%\bibitem{Fack1982} T.~Fack.
%\newblock Sur la notion de valeur caract\'eristique.
%\newblock {\it J. Operator Theory} {\bf 7} (1982), no. 2, 307--333. 
%
%\bibitem{GohbergKrein} I.~Gohberg and M.~Kre\u{i}n.
%\newblock {\it Introduction to the theory of linear nonselfadjoint operators.}
%\newblock Translations of Mathematical Monographs, Vol. 18 American Mathematical Society, Providence, R.I. 1969 xv+378 pp. 

%\bibitem{GVF} J. Gracia-Bondía, J. V\'{a}rilly and H. Figueroa.
%\newblock {\it Elements of noncommutative geometry.}
%\newblock Birkhäuser Advanced Texts: Basler Lehrbücher. Birkhäuser Boston, Inc., Boston, MA, 2001. xviii+685 pp.
%
%\bibitem{Hall_2013} B.~Hall.
%\newblock {\it Quantum theory for mathematicians.}
%\newblock Graduate Texts in Mathematics, 267. Springer, New York, 2013. xvi+554 pp.
%
\bibitem{Herb} Herbst, I. 
\newblock Spectral and scattering theory for Schr\"odinger operators with potentials independent of $|x|.$
\newblock  Am. J. Math. 113(3), 509–565 (1991)

\bibitem{HSk1} Herbst, I., Skibsted, E.
\newblock  Quantum scattering for potentials homogeneous of degree zero. 
\newblock In: Mathematical Results in Quantum Mechanics (Taxco, 2001), Contemp. Math., vol. 307, pp. 163–169. American
 Mathematical Society, Providence (2002)

\bibitem{HSk2} Herbst,I., Skibsted,E.
\newblock Quantum scattering for potentials independent of $|x|:$
asymptotic completeness for high and low energies.
\newblock Commun. Partial Differ. Equ. 29(3–4), 547–610 (2004)

%\bibitem{HoerI} L.\, H\"ormander.
%\newblock {\it The Analysis of Linear Partial Differential Operators I.}
%\newblock Grundlehren der Mathematischen Wissenschaften, 256. {\it Springer-Verlag, Berlin,} 1983. ix+391 pp.
%\bibitem{HoerIII} L.\, H\"ormander.
%\newblock {\it The Analysis of Linear Partial Differential Operators III.}
%\newblock Grundlehren der Mathematischen Wissenschaften, 274. {\it Springer-Verlag, Berlin,} 1985. viii+525 pp.

%\bibitem{KatoBook} T.\,Kato.
%\newblock {\it Perturbation theory for linear operators.}
%\newblock Springer-Verlag, Berlin, Heidelberg, New York, 1980
%
%\bibitem{Lang_genetic} R.~Lang.
%\newblock {\it Spectral theory of random Schrödinger operators. A genetic introduction.}
%\newblock Lecture Notes in Mathematics, 1498. Springer--Verlag, Berlin, 1991. viii+125 pp. 
%
%\bibitem{LeSZ_cwikel} G.~Levitina, F.~Sukochev and D.~Zanin.
%\newblock Cwikel estimates revisited.
%\newblock {\it Proc. Lond. Math. Soc.} (2019) (in press).

%\bibitem{LSZ2012} S. Lord, F. Sukochev and D. Zanin.
%\newblock {\it Singular traces: theory and applications}.
%\newblock Walter de Gruyter, Vol. 46, 2012.

%\bibitem{PasFig} L.~Pastur and A.~Figotin.
%\newblock {\it Spectra of random and almost-periodic operators.}
%\newblock Grundlehren der Mathematischen Wissenschaften, 297. {\it Springer-Verlag, Berlin,} 1992. viii+587 pp.
%
%\bibitem{PS_acta} D.~Potapov and F.~Sukochev.
%\newblock Operator-Lipschitz functions in Schatten-von Neumann classes.
%\newblock {\it Acta Math.} {\bf 207} (2011), no. 2, 375–-389. 
%
%\bibitem{ReedSimonII} M.~Reed and B.~Simon.
%\newblock {\it Methods of modern mathematical physics. II. Fourier analysis, self-adjointness.}
%\newblock Academic Press [Harcourt Brace Jovanovich, Publishers], New York-London, 1975. xv+361 pp. 
%
\bibitem{ReedSimonIV} M.~Reed and B.~Simon.
\newblock {\it Methods of modern mathematical physics. IV. Analysis of operators.}
\newblock Academic Press [Harcourt Brace Jovanovich, Publishers], New York-London, 1978. xv+396 pp. 
%
%\bibitem{Rudin_functional} W.~Rudin.
%\newblock {\it Functional analysis.}
%\newblock Second edition. International Series in Pure and Applied Mathematics. McGraw-Hill, Inc., New York, 1991. xviii+424 pp.
%
\bibitem{Ruelle} D.~Ruelle, 
\newblock {\it Statistical mechanics: rigorous results.}
\newblock W. A. Benjamin, Inc. 1969

%\bibitem{SS2013} A.~Sedaev and F.~Sukochev.
%\newblock Dixmier measurability in Marcinkiewicz spaces and applications.
%\newblock {\it J. Funct. Anal.} 265 (2013), no. 12, 3053--3066. 
%
%\bibitem{SSUZ2015} E.~Semenov, F.~Sukochev, A.~Usachev and D.~Zanin.
%\newblock Banach limits and traces on $\clL_{1,\infty}$.
%\newblock {\it Adv. Math.} 285 (2015), 568--628. 
%
\bibitem{Shubin1979} M.~Shubin.
\newblock {\it The spectral theory and the index of elliptic operators with almost periodic coefficients.}
\newblock Uspekhi Mat. Nauk, Volume 34, Issue 2(206), 95--135.
%
% \bibitem{Simon1979} B.~Simon.
% \newblock {\it Trace ideals and their applications.}
% \newblock Second edition. Mathematical Surveys and Monographs, 120. American Mathematical Society, Providence, RI, 2005. viii+150 pp. 

\bibitem{SimonSemigroups} B.~Simon.
\newblock Schr\"odinger semigroups.
\newblock {\it Bull. Amer. Math. Soc.} (N.S.) 7 (1982), no. 3, 447--526. 

\bibitem{simon_functional} B.~Simon.
\newblock {\it Functional integration and quantum physics.}
\newblock Second edition. AMS Chelsea Publishing, Providence, RI, 2005. xiv+306 pp. 

\bibitem{Zw} M.~Zworski.
\newblock {\it Semiclassical analysis.}
\newblock AMS, Graduate Studies in Mathematics, vol. 138

\end{thebibliography}

\documentclass{amsart}
\usepackage{amssymb,graphicx}
\usepackage{mathrsfs}
\usepackage{color}
\usepackage{enumerate}
\usepackage{a4wide}
\usepackage[colorlinks=true,citecolor=blue,linkcolor=blue]{hyperref}

\newcommand{\Rl}{\mathbb{R}}
\newcommand{\mbR}{\mathbb{R}}

\newcommand{\Lc}{\mathcal{L}}

\newcommand{\Tr}{\mathrm{Tr}}

\newcommand{\nexp}[1]{\exp(#1)}

\def\XXint#1#2#3{{\setbox0=\hbox{$#1{#2#3}{\int}$ }
\vcenter{\hbox{$#2#3$ }}\kern-.6\wd0}}

\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{remark}

\title{The density of states depends on the domain}
\author{N.\,Azamov, E.\,McDonald, D.\,Zanin and F.\,Sukochev}

\date{\today}

\begin{document}

\begin{abstract} In this short note we demonstrate that the definition of the density of states of a Schr\"odinger operator with bounded potential in general depends on the choice of the domain undergoing the thermodynamic limit. 
\end{abstract}
\maketitle{}

\section{Introduction}

The density of states of a Schr\"odinger operator is defined as the averaged number of states per unit of volume and is a quantity of great importance in condensed matter physics. The Laplace transform of the density of states is the partition function which often serves as the initial point of theoretical investigation of a statistical mechanical system, see e.g. \cite{Ruelle}. Its definition involves restriction of the system to a sequence of finite volume domains $\{\Omega_n\}_{n=0}^\infty$ such that $\Omega_n\to \Rl^d$ and taking the thermodynamic limit $n\to\infty.$ This process involves ambiguities in the choice of boundary conditions and of the shape of the domains $\{\Omega_n\}_{n=0}^\infty.$ It is well-known that the sequence $\{\Omega_n\}_{n=0}^\infty$ must at least satisfy some F\o{}lner condition for the limit to be meaningful.

It has been known for some time that in general the density of states is independent of the boundary conditions \cite{DIT2001}, \cite[Theorem C.7.4]{SimonSemigroups}. 
Most of the literature is devoted to potentials which are periodic or almost periodic, or random potentials satisfying some ergodicity condition. In these cases it is usually the case that the choice of F\o{}lner sequence $\{\Omega_n\}_{n=0}^\infty$ is irrelevant \cite{AiWa,Beresin_Shubin,CarmonaLacroix}. The result of this note is that in general the choice of approximating F\o{}lner sequence does matter.

We prove that there exist potentials $V\in L_{\infty}(\Rl^d)$
such that the density of states defined with the sequence of balls $\{B(0,n)\}_{n=0}^\infty$ differs from that defined by the sequence of boxes $\{[-n,n]^d\}_{n=0}^\infty$. 
The potentials which we use for this purpose are radially homogeneous. That is, the ones which obey the condition $V(tx) = V(x),$ $x \in \mbR^d,  t > 0.$ Besides our paper \cite{AMSZ}, a handful of articles concerning Schr\"odinger operators with radially homogeneous potentials exist \cite{Herb,HSk1,HSk2}. The idea of the proof is that the thermodynamic limit $\Omega_n \to \mbR^d$ for radially homogeneous potentials corresponds to the semi-classical limit $\hbar \to 0.$ 


\section{Proofs}
For a bounded open subset $0 \in \Omega$ of $\Rl^d$, we denote by $\Delta_{\Omega}$ the Laplace operator in~$\Omega$ with Dirichlet boundary conditions, and by $|\Omega|$ we denote the Lebesgue measure of $\Omega$. Let $L_2(\Omega)$ denote
the Hilbert space of almost-everywhere equivalence classes of square integrable functions on $\Omega$ with the Lebesgue measure. Denote by $\Tr_{L_2(\Omega)}$ the operator trace on the ideal $\Lc_1(L_2(\Omega))$ of trace class
operators on $L_2(\Omega).$
Let~$V$ be a bounded measurable real-valued function on $\Rl^d$. We denote by $M_V$ the operator on $L_2(\Rl^d)$ of pointwise multiplication by $V$.
It is known that if for all $t > 0$ there exists the limit
\begin{equation} \label{F: thermo lim = DOS}
     \lim_{R\to\infty} \frac{1}{|R\Omega|}\Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_V)})
\end{equation}
then there exists a unique measure $\nu_{V,\Omega}$ on $\Rl$ such that
\begin{equation*}
    \lim_{R\to\infty} \frac{1}{|R\Omega|}\Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_V)}) = \int_{\Rl} \nexp{-t\lambda} \,d\nu_{V,\Omega}(\lambda),\quad t > 0.
\end{equation*}
See \cite[Proposition C.7.2]{SimonSemigroups}. The measure $\nu_{V,\Omega}$ is called the \emph{density of states}.

\begin{theorem}\label{main_result}
    Let $V \in L_\infty(\Rl^d)$ be a radially homogeneous potential, that is, for all $t > 0$ and for all $x\in \Rl^d$ we have $V(tx) = V(x)$. Then
        for every bounded open set $\Omega$ containing zero and with piecewise smooth boundary, the density of states $\nu_{V,\Omega}$ exists and is given by the formula
        \begin{equation*}
             \int_{\Rl} \nexp{-t\lambda} \,d\nu_{V,\Omega}(\lambda) = (4\pi t)^{-\frac{d}{2}}\frac{1}{|\Omega|}\int_{\Omega} \nexp{-tV(x)}\,dx,\quad t > 0.
        \end{equation*}
\end{theorem}
The above formula for the Laplace transform of $\nu_{V,\Omega}$ yields the following for the integrated density of states function $\lambda\mapsto\nu_{V,\Omega}(-\infty,\lambda],$
\begin{equation*}
    \nu_{V,\Omega}(-\infty,\lambda] = \frac{1}{|\partial \Omega|}\int_{\partial \Omega} \nu_{V(\sigma),\Omega}(-\infty,\lambda]\,d\sigma.
\end{equation*}
Here, $|\partial \Omega|$ is the $(d-1)$-Hausdorff measure of $\partial \Omega$ and $d\sigma$ is the corresponding measure, and $V(\sigma)$ is the value of $V$ at the point $\sigma\in \partial \Omega.$
This is a generalisation of the formula in \cite[Theorem 2.1]{AMSZ} to arbitrary domains, but Theorem \ref{main_result} is stronger in that it also proves existence of the density of states.

We give a proof of Theorem \ref{main_result} based on the following well-known semiclassical Weyl law, see e.g. \cite[Theorem C.6.2]{SimonSemigroups},
\cite[Theorem 10.1]{simon_functional}. For the special case of smooth $V$, see \cite[Chapter 6]{Zw}.
\begin{theorem}\label{semiclassical_weyl_law}
    Let $\Omega\subset \Rl^d$ be a bounded open set with piecewise smooth boundary, and let $V \in L_{\infty}(\Omega)$. Then
    \begin{equation*}
        \lim_{\hbar\to 0} \hbar^d\Tr_{L_2(\Omega)}(\nexp{-t(-\hbar^2\Delta_{\Omega}+M_V)}) = (4\pi t)^{-\frac{d}{2}}\int_{\Omega} \nexp{-tV(x)}\,dx.
    \end{equation*}
\end{theorem}

The following lemma is a key observation.
\begin{lemma}\label{rescaling_lemma}
    Let $R > 0$, and let $0 \in \Omega\subset \Rl^d$ be an open set with piecewise smooth boundary. If $V \in L_{\infty}(\Rl^d)$ is homogeneous, then
    \begin{equation*}
        \Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_{V})}) = \Tr_{L_2(\Omega)}(\nexp{-t(-R^{-2}\Delta_{\Omega}+M_V)}).
    \end{equation*}
\end{lemma}
\begin{proof}
    Let $U_R$ denote the unitary mapping
        $U_R:L_2(R\Omega)\to L_2(\Omega)$
    given by
    \begin{equation*}
        (U_Ru)(x) = R^{\frac{d}{2}}u(Rx),\quad x \in \Omega.
    \end{equation*}
    For any operator $T \in \Lc_1(L_2(R\Omega))$, we have
    \begin{equation}\label{unitary_trace_conjugation}
      \Tr_{L_2(R\Omega)}(T) = \Tr_{L_2(\Omega)}(U_RTU_R^*). 
    \end{equation}
    Since  $ U_RM_VU_R^* u(x) = V(Rx)u(x)$ and since $V$ is radially homogeneous
     we have
      $ M_V = U_RM_VU_R^*$
    (on the left hand side, $M_V$ is understood as acting on $L_2(\Omega)$ and on the right it acts on $L_2(R\Omega)$).
Combining this with 
      $  U_R\Delta_{R\Omega}U_R^* 
           = R^{-2}\Delta_{\Omega},$
we conclude that 
    \begin{align*}
        \Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_{V})}) &\stackrel{\text{\eqref{unitary_trace_conjugation}}}{=} \Tr_{L_2(\Omega)}(U_R\nexp{-t(-\Delta_{R\Omega}+M_{V})}U_R^*)\\
                                                               &= \Tr_{L_2(\Omega)}(\nexp{-t(-R^{-2}\Delta_{\Omega}+M_V)}).
    \end{align*}
\end{proof}

\begin{proof}[Proof of Theorem \ref{main_result}]
    By Lemma \ref{rescaling_lemma}, we have
    \begin{equation*}
        \frac{1}{|R\Omega|}\Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_V)}) = \frac{1}{|\Omega|}R^{-d}\Tr_{L_2(\Omega)}(\nexp{-t(-R^{-2}\Delta_{\Omega}+M_V)})
    \end{equation*}
    Let $\hbar = R^{-1}$, so that
    \begin{equation*}
        \frac{1}{|R\Omega|}\Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_V)}) = \frac{1}{|\Omega|} \hbar^d \Tr_{L_2(\Omega)}(\nexp{-t(-\hbar^2\Delta_{\Omega}+M_V)}).
    \end{equation*}
    According to Theorem \ref{semiclassical_weyl_law}, the limit as $\hbar\to 0$ (equivalently, as $R\to\infty$) exists, and 
    \begin{align*}
        \lim_{R\to\infty} \frac{1}{|R\Omega|}\Tr_{L_2(R\Omega)}(\nexp{-t(-\Delta_{R\Omega}+M_V)}) &= \frac{1}{|\Omega|}\lim_{\hbar\to 0}\hbar^d \Tr_{L_2(\Omega)}(\nexp{-t(-\hbar^2\Delta_{\Omega}+M_V)})\\
                                                                                               &= \frac{1}{|\Omega|}(4\pi t)^{-\frac{d}{2}}\int_{\Omega} \nexp{-tV(x)}\,dx.
    \end{align*}
    Hence in this case the limit in \eqref{F: thermo lim = DOS} exists, and we deduce the existence of a density of states measure $\nu_{V,\Omega}.$ The above computation yields the formula
    \begin{equation*}
        \int_{\Rl} \nexp{-t\lambda}\,d\nu_{V,\Omega}(\lambda) = \frac{1}{|\Omega|}(4\pi t)^{-\frac{d}{2}}\int_{\Omega} \nexp{-tV(x)}\,dx,\quad t > 0. 
    \end{equation*}
\end{proof}

\begin{theorem} 
There exist radially homogeneous potentials $V\in L_{\infty}(\Rl^d)$ for which 
          $  \nu_{V,[-1,1]^d} \neq \nu_{V,B(0,1)},  $
        where $B(0,1)$ is the open ball of radius $1$ in $\Rl^d.$ 
\end{theorem}
\begin{proof} 
    It follows from Theorem \ref{main_result} that for any bounded open set $\Omega$ containing zero and with piecewise smooth boundary we have
    \begin{equation*}
        \int_{-\infty}^\infty \nexp{-t\lambda} \,d\nu_{V,\Omega}(\lambda) = (4\pi t)^{-\frac{d}{2}}  \left(1+\frac{t}{|\Omega|}\int_{\Omega} V(x)\,dx+O(t^2)  \right),\quad t\to 0.
    \end{equation*}
    Hence, to prove the claim it suffices to give an example of a radially homogeneous potential $V \in L_{\infty}(\Rl^d)$ such that
    \begin{equation*}
        \frac{1}{|[-1,1]^d|}\int_{[-1,1]^d} V(x)\,dx \neq \frac{1}{|B(0,1)|} \int_{B(0,1)} V(x)\,dx.
    \end{equation*}
    Such an example is given by the function
    \begin{equation*}
       V(x) = \frac{|x_1x_2|}{|x_1|^2+|x_2|^2},\quad x = (x_1,\ldots,x_d) \in \Rl^d.
    \end{equation*}
    Indeed, for $d=2$ we may compute
    \begin{equation*}
        \frac{1}{|[-1,1]^2|}\int_{[-1,1]^2} V(x)\,dx = \int_0^1\int_0^1 \frac{x_1x_2}{x_1^2+x_2^2}\,dx_1dx_2 = \frac{1}{2}\log(2)
    \end{equation*}
    and
    \begin{equation*}
        \frac{1}{|B(0,1)|}\int_{B(0,1)} V(x)\,dx = \frac{4}{\pi} \int_0^{\frac{\pi}{2}}\int_0^1 r\cos(\theta)\sin(\theta)\, drd\theta = \frac{1}{\pi}.
    \end{equation*}
    For $d>2$ the computation is identical.
\end{proof}

\end{document}